IFMIF/EVEDA Project Committee #30

20 & 21 October 2022

The 30th Project Committee (PC#30) was held on the 20th and 21st of October 2022 at the QST Rokkasho Fusion Institute. For the first time since the Covid outbreak, it took place not only by video conference, but also in-person, welcoming members and experts from Japan and Europe.

The first day was dedicated to the update of the project status and the outlook for 2023 with many presentations.

The main topics highlighted since last spring’s PC#29 were the following:

  • The injector continuous wave campaign with the 12 mm plasma electrode reached 160 mA extracted current. The requested beam characteristics have not yet reached the requirements and further operation is needed;
  • The continuous wave conditioning of the RFQ was suspended to test the RF power couplers with brazed-window, as well as to improve the design of the O-ring couplers;
  • The SRF Linac assembly started in the Joint Research Building at Rokkasho on 22 August 2022, following the completion of the ultrasonic cleaning and high-pressure rinsing of the superconducting solenoids;
  • The Fusion Neutron Source Design (FNSD) activities made significant progress both on the Lithium Facility and on the Fusion Neutron Sources, with many technical exchanges illustrating a strong collaboration between Japan EU-HT and JA-HT.
RF power couplers presentation and remote PC members
PC#30 participants both in person and remotely

The second day was dedicated to the site tour with the visit of:

  • the accelerator vault (see below picture),
  • the RF area, to have a look at the High-Power Test Bench among many systems,
  • the Joint Research building, to see the progress made in the SRF Linac assembly, and
  • the Central Control Room.

On the afternoon, the PC members reviewed and approved the Record of Conclusion. Hopefully, the next Project Committee, planned on March 16 and 17, 2023, will be only face-to-face.

Site tour – visit of the Accelerator vault by PC members and experts

Weekly LIPAc highlights (23/May-27/May)

Injector operation

The measurements with the new EMU (Emittance Meter Unit) were interrupted at the end of April because of an HV standoff issue due to a broken ceramic insulator placed at the repeller electrode. As the replacement of the ceramic insulator implies to open the injector chamber and break the vacuum, it was pointless to resume EMU qualitative characterization. It was then decided to change the PE (plasma electrode) to the 12mm size (as already planned).

During the inspection the BN (Boron Nitride) disks were also found strongly damaged and therefore replaced, it was to be expected considering the number of beam extraction and operational hours since their last replacement. We have currently a major intervention on going on the RFQ couplers. The vault is then open in the day time and injector operation can resume at around 17h30, for a shift until 21h30. While during the night and over weekends, plasma conditioning is performed. Until now efforts have been made to increase the extracted beam current remaining at 5 or 10%DC (duty cycle); current was progressively increased from 130 to 150 mA before some arcs appeared. From this week, beam characterization with the 12mm PE will start.

RFQ couplers

The RFQ cavity is currently vented with N2 at atmospheric pressure. The intervention on the couplers was slowed down mainly due to the technical difficulties (including long and delicate preparation work) and necessity to train adequately the staff by the experts. INFN supported remote training sessions on a spare coupler before the intervention was authorized. After the previous leak tests, the RFQ coupler#3b was suspected to leak and thus it has been the first disassembled and inspected. The window was found with sign of Cu deposition and with partly melted o-ring. The RF window parameters have been measured with network analyzer and the data are analyzed. Considering the time needed to inspect each coupler and replace the o-rings, we will not be able to have few weeks of beam operation in July. Accelerating the activity on the couplers implies unacceptable and not justified technical risks, additionally contractual maintenance onsite starts on August 1st.

We have also started to prepare a high power test bench (using a bridge cavity that was available onsite) to be able to tests at high power the new RF windows before their installation on the RFQ itself. This will eliminate the risk that a window with imperceptible impurities could break when used at full RF power. The bridge cavity is currently being characterized that is the measurement of RF parameters and tuning, while mechanical interfaces, possible location, cooling options are studied.

RF system

After receiving the spares onsite, the circulator of the RF-RFQ chain#1b was repaired following the instructions provided by the supplier. We are ready to perform some quick functional tests in these days. As the possibility to have beam operation is summer has been discarded, the circulator will be finally shipped back to the supplier for a repair at their premises. Even though it wastechnically the best option since the beginning, but it was not adopted for the long time needed (4 months), non-compatibility with operation in summer and the non-null chances of success of the repair onsite. As the conditioning will resume now in October, it becomes the appropriate option.

We target to have the #1b chain operational by end of October along with the RFQ. Depending on the outcome of inspections and analyses of the RF windows, the best option for the experimental campaign after the summer break will be considered: resume RFQ conditioning aiming at CW or resume beam operation at low power, low DC.

Since beginning of May, after more than 2 years of entry ban in Japan, we could welcome 5 new colleagues from Spain. 2 senior experts from CIEMAT (and old acquaintance of the project) and 3 new engineers for Granada University. We could immediately feel the positive energy and the increase of our working force in the commissioning of the HEBT and BD and in the maintenance activities. Among whom, 4 of them will stay in Rokkasho for at least 1 year.

RFQ coupler#3b during the inspection under clean booth near to the RFQ in the vault ©2022 IFMIF/EVEDA project

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (28/March-1/April)

Injector conditioning campaign

After several long runs (each one of ~10hours of CW operation) the Injector CW campaign with the 11mm PE is now considered completed. The last few days of operation were dedicated to collect data for the beam profile measurements and their analyses are on-going (see below profiles). As the RFQ conditioning campaign needs to be extended, the way forward for the injector CW campaign was agreed with CEA as follows. At first the validation of the new EMU will take place with its calibration using the 11mm PE. Then, following a short maintenance of the source (also changing the BN disk), the campaign will proceed with the 12mm PE with the goal to repeat the CW operation. The target in any case is to complete this CW campaign by mid-June, when RF-RFQ conditioning should also be completed.

Extracted current=142-145 mA, Intermediate Electrode Voltage=18 kV

Beam profile measured from solenoid magnet SOL1 @ 250A to 370A and from SOL1 @450A to 600A for 5%DC, 50%DC, and CW

RF-RFQ conditioning campaign

While the RF-RFQ conditioning reached 15%DC at nominal (for deuteron acceleration) cavity voltage of 132kV and 35%DC at 80% level cavity voltage of 106kV, it was stopped on March 30 due to vacuum interlock events. Preliminary investigation pointed out that the time needed by the system to recover the RFQ base vacuum after the vacuum peaks follows the same trend than the temperatures measured in correspondence of the RFQ couplers ceramic windows (see below graphs). Consequently, leaks at the RF couplers viton O-rings are suspected. Intensive leak checks have been performed, first by using Kr gas sprayed in pockets around the windows to be detected by the RFQ RGAs and then by using Helium, following a more standard He leak test by isolating the cryopump and using a mobile trolley. Following the increase of DC during the last months of conditioning campaign, the ceramic windows may also have experienced metallization. Further inspection of the couplers will tell us more if changing ceramic windows and O-rings should be considered. Even though spares are available on-site, the diagnosis of the issue (with the great support of INFN) is our first focus to clearly identify the coupler to be opened because the intervention will be quite invasive (with the need to disassemble the coaxial lines and disconnect them in the pit for couplers located on the north side).

Vacuum and couplers temperature trends observed during the vacuum events in the RFQ

HEBT/BD commissioning

The remote sessions with CIEMAT are progressing well. In particular the new radiation resistant RGA installed in the HEBT is now fully operative (see the below screenshot of its OPI). The noise issues at the HAL02 accelerometer in the BD was solved and the final debugging of the mass flow controllers and cRIO is on-going. The monitoring of the BPMs is ongoing after the new PV gateway service through the Data-diode was started. It appears that an overrun memory issue was causing the segmentation fault observed in the last BPMs operations. These activities will speed up when two CIEMAT colleagues together with the 3 new colleagues from Granada University will join us on-site as soon as May.

Control System

Meanwhile, the Control system supported the modification of the OPI for the isolation valve of the injector vacuum sector 2, while the modification of the chopper interlock module is on-going.

HEBT RGA operational with new OPI

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (7/March-11/March)

In the first months of 2022 the operations on-site are focused on the CW conditioning campaigns for the injector and the RFQ.

The injector conditioning campaign makes steady and quick progresses: the commissioning with 11mm plasma electrode reached 155mA of total extracted beam current in CW already several times. Fine tuning of several parameters is on-going to improve the stability of the beam (decrease the sparks, current fluctuations, improve emittance etc.) and determine a stable working point for long term CW operation.

The RF-RFQ conditioning suffered a stop in February because of a serious damage at the circulator of the RF-RFQ chain#1b. After abnormal operations and high number of arcs during conditioning in the chain#1b, the circulator was opened and the connection point between the inner plate of the circulator and the matching section of coax line was found to be broken. The EU and QST teams on-site have reacted calmly and with great spirit of collaboration and proceeded to: analyse the accident, determine the possible causes, inspect carefully all other circulators connections and define an action plan with associated risk analyses. The supplier of the circulator was contacted and we are in the process to order spare parts to repair the circulator on-site. In parallel the procurements of spares circulators will be launched. A study has been performed and associated documents and procedures have been approved to resume conditioning of the RFQ by using 7 RF-RFQ chains (out of the 8 available, i.e. with chain#1b being off). In these days a dummy load has been installed on the coaxial line of chain#1b to manage the reflected power during RF injection and RF-RFQ conditioning has just resumed.

It is difficult to have a clear idea about the impact on the schedule: the repair of the circulator on-site is planned in April and assuming that RF-RFQ conditioning will be resumed with 8 chains, we might reach CW operation by beginning of June. This will leave a couple of months for beam operation before the yearly summer maintenance break.

We had also several remote sessions with CIEMAT support to solve punctual issues (like a problem on a limit switch of a motor driven drain valve for the Beam Dump cartridge in the vault, to install the SW for a newly installed radiation resistant RGA in the HEBT) and to keep progressing in the finalization of the vacuum machine state for the HEBT.

The Diagnostic team has been working to recover full operability of the Beam Position Monitors and is currently also checking the SEM grids. The Fluorescence Profile Monitor installed in the HEBT have been moved out of the vault for a necessary recalibration that is ongoing.

We had support of a subcontractor on-site to finalize the maintenance of cooling skid (with replacement of mechanical seals in several pumps) and some re-installation work of N2 gas system in the deaerator loop of the Beam Dump skid.

Injector operation record with stable CW beam extraction

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc. 

Weekly LIPAc highlights (24/January-28/January)

The experimental campaign at the end of 2021 was successful. The first stage of Phase B+ was completed successfully by December 20 with D+ beam transported at low duty cycle and low current up to the BD. Beam operation was carried out during the day and RF-RFQ was switched to conditioning mode during the night. The duty cycle of RF-RFQ conditioning reached CW at the cavity voltage of 101 kV at the last moment of our activity on Sunday 19 December 2021.

At first, 100% was tried at 20:19 of 19 December with 20ms pulse length and 20ms repetition period (i.e. 50Hz). One event of coupler temperature over the interlock limit of 65°C at the 2A tapered cone section occurred at 20:35, however CW was kept until around 23:00.

At around 00:30 of 20 December, 100% was reached again but was not so stable.

At 07:38 of 20 December, 99.5% (mostly CW) was reached and kept it until 08:13: this was the last trial before stopping the system. See RF-RFQ conditioning summary plot for some details.

On December 20 all the systems were progressively stopped and winter maintenance with planned interventions started. We had a specialized company onsite for maintenance on the cooling skids and replacement of mechanical seals on the pumps of RF skid, RFQ skid and injector skid. After the interventions a water leak at the backup pump of injector skid could not be solved and another maintenance is planned in February. Resins for the water quality control have also been replaced in MEBT and HEBT skids. During the first week of break an alignment survey was performed in the vault for checking the difference of LEBT, RFQ and MEBT positioning with those observed in December 2020 and March 2021: the results show some systematical changes that are being analyzed carefully.

The maintenance on the injector and RF-RFQ systems was completed on 16 and 25 January respectively. The first part of 2022 will be dedicated to the injector and RF-RFQ CW campaigns. We target for the time being end of March for completing both CW campaigns.

The injector underwent standard maintenance and the Plasma Electrode was replaced with the 11mm size as agreed. The chopper was disassembled specifically for the CW campaign. The Emittance Meter Unit was checked. After the PPS test the HV conditioning started. 100 keV D+ beam extraction from the 11 mm plasma electrode is on-going for reaching 100 % DC at current higher than 150 mA: 120 mA in 10 ms pulse width over 300 ms period was reached as the first target. A trigger signal to synchronize the acquisition camera with the source pulse was implemented.

In the RF system a new SW for the LLRF autocalibration in the master chain was installed and confirmed to be operational; also the White Rabbit Grand Master was updated to suppress the desynchronization issue. In the RFQ pick-up sensors the calibration of data acquisition was completed and the data are being checked by INFN. After inspection of the DC pickup port of SSPA-A for buncher #2 (that showed signal issues last year) a test was carried out but it is confirmed that a new pick-up should be installed. The intervention is planned next week.

In the BD skid O2 control is unavailable due to N2 gas generator failure, and steady circulation through BD cartridge is to be avoided (currently the skid is kept running bypassing the cartridge). Currently we are facing a calibration issue in the pressure sensors of the BD skid and we are investigating the issue with help of CIEMAT. Remote sessions have resumed since few weeks as Japan has renewed the entry ban and CIEMAT mission is forced to be postponed.

All Fluorescence Profile Monitors (FPMs) have been disassembled and will be recalibrated (also with remote support by CIEMAT).

Beam Position Monitors (BPMs) Clock Board issue is under investigation and procurement of spares is being launched.

In the pictures below the beam operation from Local Control Room and Central Control Room

Disclaimer: the information reported is not meant to be technically complete and doesn’t cover all the activities currently carried out on LIPAc.